OpenStudy (anonymous):

how would you graph this plz help y greater than or equal to -3x + 1 y greater than or equal to 1 over 2x + 3

OpenStudy (anonymous):

y greater than or equal to -3x + 1 \[ y \ge -3x + 1\]

OpenStudy (anonymous):

y greater than or equal to 1 over 2x + 3 \[ y \ge \frac{ 1 }{ 2x + 3 }\]